OpenStudy (anonymous):

Given that (equation below) by setting up the Riemann sums and then taking the limit.

OpenStudy (anonymous):

\[\sum_{i=1}^{n} i^3 = ((n(n+1))/2)^2, compute \int\limits\limits\limits_{0}^{1} (x^3 + 1)dx\]

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I don't get what the sigma notation part means. Other than that I can do intergration and Riemann sums fairly well.

OpenStudy (zarkon):

the sigma notation is just shorthand for adding things \[\sum_{i=1}^{n} i^3=1^3+2^3+\cdots +n^3\]

OpenStudy (anonymous):

Then the equation after it?